OpenStudy (anonymous):

Find a pair of factors for each number by using the difference of two squares. a. 45 b. 77 c. 112

OpenStudy (helder_edwin):

77=81-4=(9-2)(9+2)

OpenStudy (helder_edwin):

49=49-4=(7-2)(7+2)

OpenStudy (helder_edwin):

112=121-9=(11-3)(11+3)

OpenStudy (helder_edwin):

cool. fun.